Charles purchased a share of wells Fargo stock for $10. A year later he received a $4/share dividend and purchased another share for $20. What is his dollar-weighted return? What is time-weighted return?
Answer #1
  • Time \ weighted \ return = (\frac{Price_{1} - Price_{0} + Dividend}{Price_{0}} - 1) \times 100
  • Time \ weighted \ return = (\frac{20 - 10 + 4}{10} -1) \times 100
  • Time \ weighted \ return = (\frac{14}{10} -1) \times 100
  • Time \ weighted \ return = 40 \%

Money weighted return or the dollar weighted return is like an IRR, that means that the FV should equal the PV

Price_{0}= \frac{Price_{1} + Dividend}{(1+r)^1}

Here r = money weighted return

10 = \frac{20 + 4}{(1+r)^1}

r = 0.4 \ or \ 40 \%

As the time period is 1 year both have come out to be the same, however if the time period is more than 1 year, then these will come out to be different